# Origin of unexpected weak Gilbert damping in the LSMO/Pt bilayer system

### v2

@@ Abstract
[…]
@@ II Computational methodology
To understand the relative magnitude of damping and anti-damping torques in LSMO and LSMO/Pt systems, we compared the spin Hall angles ($\theta_{SH}$) in these two systems. The spin Hall angle is a crucial parameter that indicates the efficiency of the conversion from charge current to spin current, thereby influencing the strength of the spin pumping as well as spin-orbit torques [24]. It is important to note that recently Gupta et al. reported observation of a self-induced inverse spin Hall effect (ISHE) and spin pumping [25] within LSMO. It can be noted that similar self-induced ISE was observed by Tsukahara et al. [26] in permalloy at room temperature. This self-induced ISHE leads to the generation of a significant spin-pumping voltage within the LSMO film, indicating the presence of appreciable spin-to-charge conversion within the material. This finding is significant because it demonstrates a unique mechanism for spin current generation and spin-to-charge conversion that does not rely on spin currents generated in adjacent layers, as is typically seen in traditional measurement methods for the spin Hall effect.
@@ III Results and discussion
In order to calculate the longitudinal charge conductivity, we used a relaxation time, $\tau=10\text{fs}$. We also further investigate the significance of the $\tau$ in determining the absolute values of conductivity. Assuming that the dominant scattering mechanism is acoustic phonon scattering, the scattering rate due to the acoustic deformation potential (ADP) is expressed as [37, 38]
@@
The observed Gilbert damping in LSMO is larger than in LSMO/Pt, which can be rationalized by considering the details of the spin-orbit torque mechanisms: As mentioned before, in LSMO there is evidence of a self-induced ISHE, where the material itself generates a transverse charge current from an internal spin current without the need for an external interface. In this case, the absence of an external interface means there is no anti-damping mechanism from an external source (due to the injection of spin current from another material). The self-induced spin pumping effect powered by larger values of $|\theta_{\text{SH}}|$ in LSMO, therefore, leads to significant angular momentum dissipation within the material itself, contributing to larger effective damping via conversion of spin current to charge current through $\mathbf{J}_{c}=\theta_{SH}\left(\frac{\hbar}{2e}\right)(\mathbf{J}_{s}\times\mathbf{\sigma})$. Here $\mathbf{J}_{c}$ and $\mathbf{J}_{s}$ are respectively charge and spin current density. $\mathbf{\sigma}$ is the direction of the spin polarization [43].
@@
In the LSMO/Pt heterostructure, on the other hand, the presence of SHE guarantees the presence of ISHE as well. The ISHE contributes to spin pumping, where spin currents from LSMO are absorbed by Pt, leading to angular momentum dissipation as mentioned above which contributes to the overall damping in addition to the intrinsic damping. The SHE in Pt generates spin currents in response to an applied charge current is given by $\mathbf{J}_{s}=\theta_{SH}\left(\frac{2e}{\hbar}\right)(\mathbf{J}_{c}\times\mathbf{\sigma})$ which can flow back into LSMO and exert anti-damping torques. The presence of both mechanisms therefore compete here and balances the overall damping in the LSMO/Pt heterostructure. From the seminal work of Haney et al. [46], it can be seen that, the anti-damping torque,
$\tau_{ad}\propto\theta_{\text{SH}}\frac{\left(1-e^{-t_{HM}/\lambda_{\text{sf}}}\right)^{2}}{1+e^{-2t_{HM}/\lambda_{\text{sf}}}}$ (5)
A small value of $\theta_{\text{SH}}$ implies a less efficient spin-to-charge conversion, which directly reduces $\tau_{ad}$. However, in the experiment, an antidamping torque is still seen [7]. This can be attributed to the larger spin diffusion length present in the system. A larger spin diffusion length ($\lambda_{\text{sf}}$) means that the spin polarization decays slowly, which makes the generated spin current reach the ferromagnetic easily. The term $(1-e^{-t_{HM}/\lambda_{\text{sf}}})$ will be larger, as $e^{-t_{HM}/\lambda_{\text{sf}}}$ will decay slowly, $t_{HM}$ being the thickness of the heavy metal layers. It is to be noted that materials like permalloy have been reported with a relatively smaller value of $\theta_{\text{SH}}$ with $\lambda_{\text{sf}}$ of around 7.0 nm [47]. This indicates that while the $\theta_{\text{SH}}$ is relatively low, the $\lambda_{\text{sf}}$ is sufficiently large, allowing for effective spin transport. Also, Fache et al reported that for CoFeB/Ir heterostructure, Ir showed a non-negligible spin diffusion length of 1.3 nm in spite of the fact that $\theta_{\text{SH}}$ for the Ir is only 0.2 times that of Pt [48].
@@
To understand how the spin-diffusion length behaves in the LSMO/Pt system, we considered two monolayers of LSMO and four monolayers of Pt as shown in Fig.4. We obtained the layer-resolved spin polarization using spin-polarized DFT calculations, which provided the input data for spin polarization as a function of distance. The layer-resolved spin polarization was obtained by the equation given by [49],
$P_{L}=\frac{D_{L}^{\uparrow}\left(E_{F}\right)-D_{L}^{\downarrow}\left(E_{F}\right)}{D_{L}^{\uparrow}\left(E_{F}\right)+D_{L}^{\downarrow}\left(E_{F}\right)}$ (6)
Here L is the layer index and $D_{L}^{\uparrow(\downarrow)}\left(E_{F}\right)$ is the density of states at the Fermi energy for the majority (minority) spin electrons. By fitting these data points to an exponential decay model, $P(z)=P_{0}\exp\left(-z/\lambda_{\text{sf}}\right)+C$, with the initial spin polarization $P_{0}$ to be close to 1.0, we determined the spin diffusion length ($\lambda_{\text{sf}}$). The fitting process yielded a spin diffusion length of $\lambda_{\text{sf}}\approx 21.75\,\text{Å}$. Converting this to nanometers (nm), we found $\lambda_{\text{sf}}=2.175\,\text{nm}$. As one can see this value is quite significant, this can support the anti-damping torque as mentioned above (even with a relatively small value of $\theta_{\text{SH}}$).
Figure 4: Layer dependent spin polarization of the LSMO/Pt hetero-structure obtained using Eq.6.
In Fig.5, the partial density of states (PDOS) for the LSMO/Pt system shows that near the Fermi energy, the Pt-d bands dominate, whereas the Mn-d states—which could otherwise introduce a heavier orbital character—are found farther from $E_{F}$. This is apparent from the near-negligible Mn-d contribution around $E_{F}$ relative to the robust Pt-d weight. According to the work of Mazin et al.,[50], the significant discrepancy between DOS-based and transport-based spin polarizations generally occurs when both light s-like and heavy d-like states coexist at $E_{F}$. In contrast, here only the Pt-d orbitals are relevant near the Fermi level, suggesting that the difference between $P_{N}$ (DOS-based) and $P_{Nv^{2}}$ (transport-based) should be minimal. The predominance of a single orbital effectively bypasses the issues that occur with multiple orbital characters having varying Fermi velocities. This enhances the credibility of our density of states (DOS)-based qualitative analysis regarding the local spin asymmetry within the LSMO/Pt system.
Finally, to demonstrate that these results are an exception and not generally expected, we performed similar calculations for the Co/Pt system. The interface was considered using one monolayer of Co and one monolayer of Pt. The observed $\theta_{SH}=0.036$ for the Co/Pt bilayer system is larger than the case of a single Co-monolayer for which a quite relatively smaller value of $\theta_{SH}=0.004$ is observed. The stronger $\theta_{SH}$ in this heterostructure suggests that the Gilbert damping would increase with increasing Pt thickness due to enhanced spin pumping effect as for example reported by Mahfouzi et al. [51]. As seen from Fig.3, the Co/Pt interface has smaller longitudinal conductivity and higher SHC in comparison to the Co mono-layer something opposite in nature if one compares with LSMO and LSMO/Pt heterostructure. It should be noted here that LSMO being a half-metal is expected to have smaller intrinsic Gilbert damping as can be understood either from Kambersky’s torque-torque model [52, 53] or from the work of Gilmore et al. [54]. In this work, we have computed the intrinsic Gilbert damping within Kambersky’s torque-correlation model which is given by [55],
Where, $\hat{T}^{\pm}=\left[\hat{S}^{\pm},\hat{\mathrm{H}}_{\mathrm{SO}}\right]$ are the torque-operators. $G\left(k,E_{\mathrm{F}}\right)=[E+i\eta-H(k)]^{-}1$ is the Green’s function at Fermi energy $E_{F}$. The Hamiltonian ${H}={H}_{0}+{H}_{xc}+{H}_{SOI}$ describes the electronic structure of the system. The paramagnetic band structure is described by $H_{0}$, and ${H}_{\text{exc }}$ describes the effective local electron-electron interaction, which is treated within a spin-polarized semi-local Kohn-Sham exchange-correlation functional approach. ${H}_{\mathrm{SO}}$ is the spin-orbit Hamiltonian and $\hat{S}^{\pm}$ are the spin operators. $N$ is the number of k-points, while $M_{s}$ is the saturation moment and $\gamma$ is the gyromagnetic ratio. $\eta$ is the broadening which was set to be 10 meV in our calculation. We obtain the intrinsic damping to be 0.0014 for the LSMO, while it is 0.012 for Pt/LSMO. As LSMO is half metal, there is no spin-flip scattering for the minority spin electrons due to the reduction in the available states, which leads to fewer scattering events that contribute to energy dissipation, thereby reducing the intrinsic damping. Therefore, the appearance of a relatively larger damping in LSMO films can only be rationalized by an extrinsic source such as self-induced ISHE.

### v3

@@ Abstract
We present a first-principles and semiclassical analysis of the recent puzzling observation that a La0.7Sr0.3MnO3 (LSMO) thin film displays larger Gilbert damping than an LSMO/Pt bilayer, the opposite of what conventional spin-pumping theory predicts. Density-functional calculations combined with Wannier interpolation give an intrinsic damping of $\alpha_{\mathrm{int}}^{\mathrm{LSMO}}\!\approx\!1.4\times 10^{-3}$, confirming that the experimentally reported high damping must be extrinsic. Guided by the self-induced inverse spin Hall effect (ISHE) already demonstrated in LSMO by Gupta et al. [Phys. Rev. B 109, 014437 (2024)], we quantify how the large spin Hall angle $|\theta_{\mathrm{SH}}|\simeq 0.093$ and the small longitudinal conductivity of LSMO lead to efficient spin-to-charge conversion inside the ferromagnet, enhancing the effective damping. For the LSMO/Pt heterostructure, we find that Pt capping strongly increases the longitudinal conductivity, reducing the interfacial $|\theta_{\mathrm{SH}}|$ to 0.007 and thus suppressing spin-pumping-induced damping. By applying a Valet–Fert analysis to our layer-resolved ab-initio spin accumulation, we extract the spin-diffusion length within Pt, which, together with the calculated conductivities and spin Hall angle, yields a non-negligible anti-damping spin–orbit torque coefficient. This coefficient indicates the potential for an antidamping spin-orbit torque (SOT) to reduce the net damping in current-biased LSMO/Pt. The observed reduction in effective damping of the LSMO/Pt bilayer thus arises from the competition between the intrinsic Gilbert damping of the LSMO film and an anti-damping torque that becomes prominent when spin pumping is suppressed by the small interfacial spin Hall angle. Our results reconcile the apparently anomalous damping hierarchy without invoking any new interfacial mechanisms, and provide a quantitative design rule via the tunable ratio of Hall to longitudinal conductivities and spin-diffusion length for engineering magnetically robust, low-loss manganite heterostructures for spin-orbitronic applications.
@@ II Computational methodology
II.1 First-Principles DFT Calculations
II.2 Wannier Interpolation and Transport Properties
To understand the relative magnitude of damping and anti-damping torques in LSMO and LSMO/Pt systems, we compared the spin Hall angles ($\theta_{SH}$) in these two systems. The spin Hall angle is a crucial parameter that indicates the efficiency of the conversion from charge current to spin current, thereby influencing the strength of the spin pumping as well as spin-orbit torques [28, 29, 30, 31]. This finding is significant because it demonstrates a unique mechanism for spin current generation and spin-to-charge conversion that does not rely on spin currents generated in adjacent layers, as is typically seen in traditional measurement methods for the spin Hall effect.
@@ II.2 Wannier Interpolation and Transport Properties
II.3 Gilbert Damping via Torque-Correlation
The intrinsic Gilbert damping was calculated within Kambersky’s torque-correlation model which is given by [34],
Where, $\hat{T}^{\pm}=\left[\hat{S}^{\pm},\hat{\mathrm{H}}_{\mathrm{SO}}\right]$ are the torque-operators. $G\left(k,E_{\mathrm{F}}\right)=[E+i\eta-H(k)]^{-}1$ is the Green’s function at Fermi energy $E_{F}$. The Hamiltonian ${H}={H}_{0}+{H}_{xc}+{H}_{SOI}$ describes the electronic structure of the system. The paramagnetic band structure is described by $H_{0}$, and ${H}_{\text{exc }}$ describes the effective local electron-electron interaction, which is treated within a spin-polarized semi-local Kohn-Sham exchange-correlation functional approach. ${H}_{\mathrm{SO}}$ is the spin-orbit Hamiltonian and $\hat{S}^{\pm}$ are the spin operators. $N$ is the number of k-points, while $M_{s}$ is the saturation moment and $\gamma$ is the gyromagnetic ratio. $\eta$ is the broadening, which was set to be 10 meV in our calculation.
@@ III.1 Electronic Structure and Wannier Validation
III.2 Intrinsic damping
It should be noted here that LSMO being a half-metal is expected to have smaller intrinsic Gilbert damping as can be understood either from Kambersky’s torque-torque model [39, 40] or from the work of Gilmore et al. [41]. We obtain the intrinsic damping to be 0.0014 for the LSMO, while it is 0.012 for Pt/LSMO. As LSMO is half metal, there is no spin-flip scattering for the minority spin electrons due to the reduction in the available states, which leads to fewer scattering events that contribute to energy dissipation, thereby reducing the intrinsic damping. Therefore, the appearance of a relatively larger damping in LSMO films can only be rationalized by an extrinsic source such as self-induced ISHE.
III.3 Longitudinal charge ($\sigma_{xx}$) and spin Hall ($\sigma_{SHC}$) conductivities
@@
In order to calculate the longitudinal charge conductivity, we used a relaxation time, $\tau=10\text{fs}$. We also further investigate the significance of the $\tau$ in determining the absolute values of conductivity. Assuming that the dominant scattering mechanism is acoustic phonon scattering, the scattering rate due to the acoustic deformation potential (ADP) is expressed as [42, 43],
@@
III.4 Self-induced ISHE in LSMO and its impact on effective damping
@@
The observed Gilbert damping in LSMO is larger than in LSMO/Pt, which can be rationalized by considering the details of the spin-orbit torque mechanisms. It is important to note that recently Gupta et al. reported observation of a self-induced inverse spin Hall effect (ISHE) and spin pumping [48] within LSMO. It can be noted that similar self-induced ISE was observed by Tsukahara et al. [49] in permalloy at room temperature. This self-induced ISHE leads to the generation of a significant spin-pumping voltage within the LSMO film, indicating the presence of appreciable spin-to-charge conversion within the material. In the case of LSMO the absence of an external interface means there is no anti-damping mechanism from an external source (due to the injection of spin current from another material). The self-induced spin pumping effect powered by larger values of $|\theta_{\text{SH}}|$ in LSMO, therefore, leads to significant angular momentum dissipation within the material itself, contributing to larger effective damping via conversion of spin current to charge current through $\mathbf{J}_{c}=\theta_{SH}\left(\frac{\hbar}{2e}\right)(\mathbf{J}_{s}\times\mathbf{\sigma})$. Here $\mathbf{J}_{c}$ and $\mathbf{J}_{s}$ are respectively charge and spin current density. $\mathbf{\sigma}$ is the direction of the spin polarization [50].
@@
In the LSMO/Pt heterostructure, on the other hand, the presence of SHE guarantees the presence of ISHE as well. The ISHE contributes to spin pumping, where spin currents from LSMO are absorbed by Pt, leading to angular momentum dissipation as mentioned above which contributes to the overall damping in addition to the intrinsic damping. The SHE in Pt generates spin currents in response to an applied charge current is given by $\mathbf{J}_{s}=\theta_{SH}\left(\frac{2e}{\hbar}\right)(\mathbf{J}_{c}\times\mathbf{\sigma})$ which can flow back into LSMO and exert anti-damping torques. The presence of both mechanisms therefore competes here and balances the overall damping in the LSMO/Pt heterostructure. A small value of $\theta_{\text{SH}}$ implies a less efficient spin-to-charge conversion. However, in the experiment, an antidamping torque is still seen [7]. This can be attributed to the larger spin diffusion length present in the system.
The generation of spin-orbit torques, such as the antidamping torque, in heavy metal (HM)/ferromagnet (FM) bilayers is a complex interplay of several material and interface parameters. In the following we analyse the interface properties.
III.5 Spin-diffusion length in Pt and antidamping SOT in LSMO/Pt
To understand the behavior of spin transport and determine the spin-diffusion length ($\lambda_{\text{sf}}$) in the La0.7Sr0.3MnO3/Pt (LSMO/Pt) system, we performed first-principles calculations on a model heterostructure. This structure consisted of two monolayers (ML) of LSMO and four monolayers of Pt ($t_{\text{Pt}}\approx 0.8$ nm), as depicted in Fig. 4 . We obtained the layer-resolved spin polarization within the Pt layers using spin-polarized Density Functional Theory (DFT). The spin polarization $P(z)$ at a given position $z$ along the film orientation (normal to the interface) was calculated from the spin-dependent density of states (DOS) at the Fermi energy ($E_{F}$) using the established relation [55]:
$P(z)=\frac{D^{\uparrow}\left(E_{F},z\right)-D^{\downarrow}\left(E_{F},z\right)}{D^{\uparrow}\left(E_{F},z\right)+D^{\downarrow}\left(E_{F},z\right)},$ (6)
where $D^{\uparrow(\downarrow)}\left(E_{F},z\right)$ represents the DOS for majority (minority) spin electrons at $E_{F}$ in layer $z$.
In Fig.3, the partial density of states (PDOS) for the LSMO/Pt system shows that near the Fermi energy, the Pt-d bands dominate, whereas the Mn-d states—which could otherwise introduce a heavier orbital character—are found farther from $E_{F}$. This is apparent from the near-negligible Mn-d contribution around $E_{F}$ relative to the robust Pt-d weight. According to the work of Mazin et al.,[56], the significant discrepancy between DOS-based and transport-based spin polarizations generally occurs when both light s-like and heavy d-like states coexist at $E_{F}$. In contrast, here only the Pt-d orbitals are relevant near the Fermi level, suggesting that the difference between $P_{N}$ (DOS-based) and $P_{Nv^{2}}$ (transport-based) should be minimal. The predominance of a single orbital effectively bypasses the issues that occur with multiple orbital characters having varying Fermi velocities. This ensures the credibility of our density of states (DOS)-based qualitative analysis regarding the local spin asymmetry within the LSMO/Pt system in a reasonable manner.
Within the framework of the diffusive Valet-Fert theory [57], a stationary spin accumulation $\mu_{s}(z)=\mu_{\uparrow}(z)-\mu_{\downarrow}(z)$ within a non-magnetic material obeys the diffusion equation:
$\frac{d^{2}\mu_{s}(z)}{dz^{2}}=\frac{\mu_{s}(z)}{\lambda_{\mathrm{sf}}^{2}}.$ (7)
In the linear-response regime, where the induced spin accumulation is small, it is reasonable to assume that the DFT-calculated spin polarization $P(z)$ is proportional to $\mu_{s}(z)$. The general solution to Eq. (7) for a film of thickness $t_{\text{Pt}}$, with appropriate boundary conditions (e.g., $z=0$ at the LSMO/Pt interface and $z=t_{\text{Pt}}$ at the Pt/vacuum interface), can be expressed in a hyperbolic cosine form [57, 58],
$P(z)=P_{0}\frac{\cosh\!\bigl{[}(t_{\mathrm{Pt}}-z)/\lambda_{\mathrm{sf}}\bigr{]}}{\cosh\!\bigl{(}t_{\mathrm{Pt}}/\lambda_{\mathrm{sf}}\bigr{)}}+C,$ (8)
where $P_{0}$ is the spin polarization at the interface ($z=0$), $\lambda_{\mathrm{sf}}$ is the spin-diffusion length, and $C$ is an offset constant. This hyperbolic form reduces to a simpler exponential decay, $P(z)\approx P_{0}\exp(-z/\lambda_{\mathrm{sf}})+C$, only in the thick-film limit where the film thickness $t_{\text{Pt}}$ is much greater than the spin-diffusion length ($t_{\text{Pt}}\gg\lambda_{\mathrm{sf}}$).
In the experimental setup by Gupta et al. [7], the Pt layer thickness was 3 nm. A direct first-principles treatment of such a thick slab would be computationally prohibitive. Therefore, our DFT calculations employed a thinner 4-monolayer Pt film ($t_{\text{Pt}}\approx 0.8$ nm). Initially, an exponential decay model was considered for fitting the DFT-derived $P(z)$ data, which yielded $\lambda_{\text{sf}}\approx 2.175$ nm. However, this result ($\lambda_{\text{sf}}\approx 2.7\times t_{\text{Pt}}$) indicated that the condition $t_{\text{Pt}}\gg\lambda_{\text{sf}}$ was not met, making the exponential approximation inappropriate for our DFT slab. Therefore, we again fitted the layer-resolved DFT polarization data from the 4 ML Pt slab directly to the more general hyperbolic cosine model given by Eq. (8). This fitting procedure yielded a spin-diffusion length of $\lambda_{\text{sf,Pt}}\approx 0.7975$ nm ($7.975$). This value, where $\lambda_{\text{sf,Pt}}\approx t_{\text{Pt}}$, confirms the necessity of using the hyperbolic model for our computationally constrained thin Pt slab. While this $\lambda_{\text{sf,Pt}}$ is shorter than some bulk Pt values, it is a plausible result for an ultrathin film where interface scattering and confinement effects can significantly influence spin transport.
Using the spin-diffusion length obtained as above and employing the semiclassical model developed by Haney et al. [59], which provides an expression for the dimensionless damping-like torque coefficient, $\tau_{ad}$ as,
$\tau_{ad}$ $=\theta_{\text{SH}}\frac{(1-e^{-t_{\text{Pt}}/\lambda_{\text{sf,Pt}}})^{2}}{1+e^{-2t_{\text{Pt}}/\lambda_{\text{sf,Pt}}}}$
×$\quad\times\frac{|\tilde{G}^{\uparrow\downarrow}|^{2}+\text{Re}[\tilde{G}^{\uparrow\downarrow}]\tanh^{2}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})}{|\tilde{G}^{\uparrow\downarrow}|^{2}+2\text{Re}[\tilde{G}^{\uparrow\downarrow}]\tanh^{2}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})+\tanh^{4}(t_{\text{Pt}}/\lambda_{\text{sf,Pt}})}$ (9)
[…]
Therefore, for the ultrathin LSMO/Pt heterostructure, the efficient utilization of the Pt layer thickness relative to its $\lambda_{\text{sf,Pt}}$ and effective interface spin transmission contribute significantly to the SOT efficiency, compensating to some extent for the reduced effective $\theta_{\text{SH}}$ of the bilayer.
Figure 4: Layer dependent spin polarization of the LSMO/Pt hetero-structure obtained by fitting the DFT-data with the Eq.8 (orange line). For completeness, an exponential fit with $P(z)\approx P_{0}\exp(-z/\lambda_{\mathrm{sf}})+C$ (blue line) is also shown.
III.6 Contrast with prototypical Co/Pt bilayer
Finally, to demonstrate that the above-mentioned results are an exception and not generally expected, we performed similar calculations for the Co/Pt system. The interface was considered using one monolayer of Co and one monolayer of Pt. The observed $\theta_{SH}=0.036$ for the Co/Pt bilayer system is larger than the case of a single Co-monolayer, for which a quite relatively smaller value of $\theta_{SH}=0.004$ is observed. The stronger $\theta_{SH}$ in this heterostructure suggests that the Gilbert damping would increase with increasing Pt thickness due to enhanced spin pumping effect, as for example reported by Mahfouzi et al. [60]. As seen from Fig.6, the Co/Pt interface has smaller longitudinal conductivity and higher SHC in comparison to the Co mono-layer something opposite in nature if one compares with LSMO and LSMO/Pt heterostructure.